Suppose you are choosing between two long distance telephone plans. Plan A has a monthly fee of $20 with a charge of $0.05 per minute for all long distances calls. Plan B had a monthly fee of $5.00 with charge of $.10 per minute fr Long distances calls.
1. For what number of minutes of long distances calls will te cost for the two plans be same? What will be the cost for each plan?

Answer by belle3833(46)

cost of plan 1: 20+0.05m (where m is the number of minutes you make long distance calls
cost of plan2: 5+0.10m
when are they equal?
when 20+0.05m=5+0.1m
solve for m:
20-5=0.1m-0.05m
15=0.05m
m=15/0.05
m=300 min
the cost of both plans at 300 min is 20+0.05 x 300=$35